# An Efficient Numerical Scheme for a Time-Fractional Burgers Equation with Caputo-Prabhakar Derivative

### v1

@@
Abstract.
This paper presents a numerical method to solve a time-fractional Burgers equation, achieving order of convergence $(2-\alpha)$ in time, here $\alpha$ represents the order of the time derivative. The fractional derivative is modeled by Caputo–Prabhakar (CP) formulation, which incorporates a kernel defined by the three-parameter Mittag-Leffler function. Finite difference methods are employed for the discretization of the derivatives. To handle the non-linear term, the Newton iteration method is used. The proposed numerical scheme is proven to be stable and convergent in the $L_{\infty}$ norm. The validity of the theory is supported by two numerical examples.
Key words and phrases:
Caputo–Prabhakar derivate, Fractional Burger equation, Finite difference scheme, Numerical approximation, Stability and convergence analysis
2020 Mathematics Subject Classification:
35R11, 65M06, 65M12
1. Introduction
@@
Unlike the classical definition of integer-order derivatives, fractional derivatives are defined in multiple ways, each suited to different contexts. Various perspectives have been presented by Caputo, Hadamard, Grünwald, Letnikov, Riemann, and Riesz, offering solutions to a wide range of real-world problems [3, 14, 20, 25, 27]. In this paper, we consider the Caputo–Prabhakar (CP) derivative, which is a generalization of the Caputo derivative obtained by introducing three parameter Mittag–Leffler function, termed as Prabhakar function, in the kernel [28]. For an absolutely continuous function $f\in\text{AC}^{1}(0,L)$ of order $\alpha\in(0,1)$, we can define CP derivative as
$^{\operatorname{CP}}_{~{}~{}0}D_{\rho,\alpha,\omega}^{\gamma}f(t)=\int_{0}^{t}(t-s)^{-\alpha}E_{\rho,1-\alpha}^{-\gamma}(\omega(t-s)^{\rho})f^{\prime}(s)ds,$ (1.1)
@@
is a three parameter Mittag-Leffler function, here $\alpha,\beta,\gamma$ are complex numbers. A Prabhakar function $E_{\alpha,\beta}^{\gamma}$ is an entire function of order $(\text{Re}(\alpha))^{-1}$ provided $\text{Re}(\alpha)>0$ and it reduces to many very well-known special functions for particular values of $\alpha,\beta,\gamma$. Subsequently, the integral (1.1) exists for $Re(\alpha)>0$.
@@
1.1. Mathematical model
We investigate the following type of fractional Burgers equation involving CP derivative in the time direction:
$\frac{{}^{\operatorname{CP}}\partial^{\alpha}u(x,t)}{\partial t^{\alpha}}+u(x,t)\frac{\partial u(x,t)}{\partial x}=\frac{\partial^{2}u(x,t)}{\partial x^{2}}+f(x,t),\quad(x,t)\in(0,L)\times(0,T),$ (1.2)
@@ 1.1. Mathematical model
${\left.\begin{array}[]{lc@{}c@{}r}u(x,0)=\phi(x),\quad x\in(0,L)\\ u(0,t)=u(L,t)=0,\quad t\in(0,T)\end{array}\right\}}$ (1.5)
For our case, $\alpha\in(0,1)$, $L$ and $T$ are positive real numbers. $f(x,t)$ is a given source term. The functions $f(x,t)$ and $\phi(x)$ are the sufficiently smooth known functions.
In [35], the authors addressed a similar problem using the Atangana–Baleanu Caputo derivative by applying a direct linearization technique, which led to a reduced convergence order in the temporal direction. In contrast, we employ the Newton iteration method to achieve a higher convergence rate and improved accuracy. It is important to note that the scheme presented here is a novel contribution, as its application to solve problem (1.2) has not been previously explored in the literature.
@@
2. Derivation and Analysis of the Numerical Scheme
We assume the grid points as $t_{k}=k\tau,~{}k=0,1,\dots,N$ in time direction and for space direction, $x_{i}=ih,~{}i=0,1,\dots,M$, where $\tau=\frac{T}{N}$ and $h=\frac{L}{M},$ for $M,N\in\mathbb{Z}^{+}$. For the discretization of time-derivative of (1.2), we use finite difference method (FDM) in the definition (1.1). So at a grid point $(x_{i},t_{k})$, we have
@@ 2. Derivation and Analysis of the Numerical Scheme
[…]
$=$ $\frac{1}{\mu}[a_{1}u_{i}^{k}-\sum_{j=1}^{k-1}(a_{k-j}-a_{k-j+1})u_{i}^{j}-a_{k}u_{i}^{0}]+r_{k},$ (2.1)
@@
$a_{n}=n^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{n})^{\rho})-(n-1)^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{n-1})^{\rho}),$ (2.2)
@@
Lemma 2.1.
[32] Let $u(x,t)\in C^{2}[0,1]$ then for any $\alpha\in(0,1)$, the truncation error $r_{k}$ occurred in approximating the CP derivative using FDM, given by
@@
$\lvert r_{k}\rvert$ $\leq C_{1}\frac{\tau^{2-\alpha}}{2}\max_{0\leq t\leq t_{k-1}}u^{\prime\prime}(x_{i},t_{k}),$ (2.3)
@@
We approximate the space derivatives of integer order from (1.2) as following,
$\frac{\partial^{2}u(x,t)}{\partial x^{2}}|_{(x_{i},t_{k})}$ $\approx\frac{u^{k}_{i+1}-2u^{k}_{i}+u^{k}_{i-1}}{h^{2}}:=\delta^{2}_{x}u_{i}^{k},$ (2.4)
$\frac{\partial u(x,t)}{\partial x}|_{(x_{i},t_{k})}$ $\approx\frac{u^{k}_{i+1}-u^{k}_{i-1}}{2h}:=\frac{1}{2h}\Delta u_{i}^{k},$ (2.5)
@@
$=\frac{1}{6h}(u_{i}^{k}\Delta u_{i}^{k}+\Delta(u_{i}^{k})^{2}).$ (2.6)
The discretizations (2.4) and (2.5) are simply central differences, while (2.6) is based on the Galerkin method using piecewise linear test functions [22, 29]. To get the fully discrete numerical scheme for the fractional Burgers equation (1.2), we substitute the approximations at point $(x_{i},t_{k})$ from (2), (2.4), and (2.6),
$\frac{1}{\mu}[a_{1}u_{i}^{k}-\sum_{j=1}^{k-1}(a_{k-j}-a_{k-j+1})u_{i}^{j}-a_{k}u_{i}^{0}]+\frac{1}{6h}(u_{i}^{k}\Delta u_{i}^{k}+\Delta(u_{i}^{k})^{2})=\delta^{2}_{x}u_{i}^{k}+f_{i}^{k},$ (2.7)
[…]
$u^{k}_{0}=u^{k}_{M}=0,~{}~{}~{}~{}u^{0}_{i}=\phi(x_{i}),$ (2.8)
2.1. Stability and Convergence
First, we introduce some notations directly inspired by [22, 29], and we recommend these papers for a detailed explanation. Define $W=\{w=(w_{0},w_{1},\dots w_{M})|w_{0}=w_{M}=0\},$ and for any two elements $v,w\in W$ define inner product as $\langle v,w\rangle_{h}=\sum_{i=1}^{M-1}hv_{i}w_{i}$, and the corresponding norms are $\|w\|_{h}=\sqrt{\langle w,w\rangle_{h}},~{}~{}~{}~{}\|w\|_{\infty}=\max_{1\leq i\leq{M-1}}|w_{i}|.$ Given the boundary conditions by (2.8), we have $\|w\|_{\infty}\leq\frac{\sqrt{L}}{2}\|\delta_{x}w\|_{h},$ where $\delta_{x}w_{i}=\frac{1}{h}(w_{i}-w_{i-i}).$
@@ 2.1. Stability and Convergence
$\langle\delta_{x}^{2}w_{1},w_{2}\rangle_{h}=-\langle\delta_{x}w_{1},\delta_{x}w_{2}\rangle_{h}~{}~{}~{}~{}\forall w_{1},w_{2}\in W,$ (2.9)
$\langle w\Delta w+\Delta(w)^{2},w\rangle_{h}=0,~{}~{}~{}~{}\forall w\in W,$ (2.10)
where $\delta_{x}^{2}$ and $\Delta$ are given by (2.4) and (2.5), respectively.
Lemma 2.2.
For any vector solution $U=(U_{0},U_{1},U_{2},\dots,U_{N})$ to (2.7)-(2.8), the following inequality is satisfied
@@
$-\frac{N^{1-\alpha}}{2}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N})^{\rho})U_{0}^{2},$ (2.11)
where $\{a_{n}|n=1,\dots,k\}$ are given by Eq. (2.2), and are decreasing function of $n$.
@@
By taking the left hand side of the inequality (2.2), we can simplify it as follows using simple well-known relations and inequalities
@@
$a_{N}=N^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N})^{\rho})-(N-1)^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N-1})^{\rho})$, we have the inequality (2.2). ∎
Following theorem proves the boundedness of the numerical solution obtained from the proposed numerical scheme (2.7) which shows the scheme (2.7) is stable.
Theorem 2.3.
Let the solution of the proposed numerical scheme (2.7) with boundary and initial conditions (2.8) is $\{u_{j}^{k}|1\leq j\leq M-1,1\leq k\leq N\}$, then we have
@@
By taking inner product of Eq. (2.7) with $u^{k}$ then by summing up from 1 to $N$, we get
@@
In the last inequality above, we used Eq.(2.2) and Eq.(2.10). By the virtue of Young’s inequality: for all $a,b\in\mathbb{R}$ there exists $\varepsilon>0$ such that $ab\leq\varepsilon a^{2}+\frac{1}{4\varepsilon}b^{2},$ we have
@@
Now, we will discuss the convergence of our scheme for which we use energy method. Assume $v(x_{i},t_{k})=v^{k}_{i}$ be the exact solution of the problem (1.2)-(1.5), and let $\epsilon_{i}^{k}=v(x_{i},t_{k})-u^{k}_{i}$ be the error between exact solution and the approximated solution, then $\epsilon_{0}^{k}=\epsilon_{M}^{k}=0$ and $\epsilon_{i}^{0}=0~{}~{}\forall~{}~{}1\leq i\leq M-1,~{}~{}1\leq k\leq N$, and
@@
$-\frac{\mu}{6h}\left[\left(v_{i}^{k}\Delta\epsilon_{i}^{k}+\Delta(\epsilon_{i}^{k}v_{i}^{k})\right)+\epsilon_{i}^{k}\Delta v_{i}^{k}+\Delta(\epsilon_{i}^{k}v_{i}^{k})\right].$ (2.12)
@@
$+\mu\left[{(\gamma_{1})}_{i}^{k}-{(\gamma_{2})}_{i}^{k}-{(\gamma_{3})}_{i}^{k}-\frac{1}{6h}\left({(\gamma_{4})}_{i}^{k}-{(\gamma_{5})}_{i}^{k}\right)\right].$ (2.13)
We can easily calculate the following results by using initial and boundary conditions ( see [29, Lemma 3.6, Lemma 3.7])
$\langle{(\gamma_{4})}_{i}^{k},\epsilon^{k}\rangle_{h}=0,~{}~{}~{}~{}|\langle{(\gamma_{5})}_{i}^{k},\epsilon^{k}\rangle_{h}|\leq 3C_{0}h\|\epsilon^{k}\|_{h}^{2},$ (2.14)
$\sum_{k=1}^{N}\tau(\|{(\gamma_{1})}_{i}^{k}\|_{h}^{2}+\|{(\gamma_{2})}_{i}^{k}\|_{h}^{2}+\|{(\gamma_{3})}_{i}^{k}\|_{h}^{2})\leq C(\tau^{2-{\alpha}}-h^{2})^{2},$ (2.15)
@@
Now, in Eq. (2.13), taking product with $h\epsilon_{i}^{k}$ and then summing up for $i$ from 1 to $M-1$,
@@
Theorem 2.4.
The finite difference scheme (2.7) is convergent in the following sense
$\tau\sum_{k=1}^{N}\|\delta_{x}\epsilon^{k}\|_{h}^{2}\leq C(\tau^{2-\alpha}-h^{2})^{2},$ (2.16)
@@
3. Numerical Examples
Before conducting numerical examples we introduce Newton’s iterative method (see [29, 31]) to deal with the non-linear term in the numerical scheme (2.7).
3.1. Newton’s iterative method
Let us assume $\psi(u)=uu_{x}$, and hence the right hand side of (2.6) can be denoted as $\psi(u)^{k}_{i}$. By rewriting (2.7), we get

### v2

@@
Abstract
This paper presents a numerical method to solve a time-fractional Burgers equation, achieving order of convergence $(2-\alpha)$ in time, here $\alpha$ represents the order of the time derivative. The fractional derivative is modeled by the Caputo–Prabhakar (CP) formulation, which incorporates a kernel defined by the three-parameter Mittag-Leffler function. Finite difference methods are employed for the discretization of the derivatives. To handle the non-linear term, the Newton iteration method is used. The proposed numerical scheme has been shown to be stable and convergent in the norm $L_{\infty}$. The validity of the theory is supported by two numerical examples.
keywords:
Caputo–Prabhakar derivative, Fractional Burger equation, Finite difference scheme, Numerical approximation, Stability and convergence analysis
pacs:
MSC Classification]35R11, 65M06, 65M12
1 Introduction
@@
Unlike the classical definition of integer-order derivatives, fractional derivatives are defined in multiple ways, each suited to different contexts. Various perspectives have been presented by Caputo, Hadamard, Grünwald, Letnikov, Riemann, and Riesz offering solutions to a wide range of real-world problems [3, 14, 21, 26, 28]. In this paper, we consider the Caputo–Prabhakar (CP) derivative which is a generalization of the Caputo derivative obtained by introducing three parameter Mittag–Leffler function, termed as Prabhakar function, in the kernel [29]. For an absolutely continuous function $f\in\text{AC}^{1}(0,L)$ of order $\alpha\in(0,1)$, we can define CP derivative as
$^{\operatorname{CP}}_{~~0}D_{\rho,\alpha,\omega}^{\gamma}f(t)=\int_{0}^{t}(t-s)^{-\alpha}E_{\rho,1-\alpha}^{-\gamma}(\omega(t-s)^{\rho})f^{\prime}(s)ds,$ (1)
@@
is a three parameter Mittag-Leffler function, here $\alpha,\beta,\gamma$ are complex numbers. A Prabhakar function $E_{\alpha,\beta}^{\gamma}$ is an entire function of order $(\text{Re}(\alpha))^{-1}$ provided $\text{Re}(\alpha)>0$ and it reduces to many very well-known special functions for particular values of $\alpha,\beta,\gamma$. Subsequently, the integral (1) exists for $Re(\alpha)>0$. For convenience, we denote the CP derivative as
$^{\operatorname{CP}}_{~~0}D_{\rho,\alpha,\omega}^{\gamma}(\cdot):=\frac{{}^{\operatorname{CP}}\partial^{\alpha}(\cdot)}{\partial t^{\alpha}}.$
@@
1.1 Mathematical model
We investigate the following fractional Burgers equation involving CP derivative in the time direction:
$\frac{{}^{\operatorname{CP}}\partial^{\alpha}u(x,t)}{\partial t^{\alpha}}+u(x,t)\frac{\partial u(x,t)}{\partial x}=\frac{\partial^{2}u(x,t)}{\partial x^{2}}+f(x,t),\quad(x,t)\in(0,L)\times(0,T),$ (2)
@@ 1.1 Mathematical model
${\left.\begin{array}[]{lc@{}c@{}r}u(x,0)=\phi(x),\quad x\in(0,L),\\ u(0,t)=u(L,t)=0,\quad t\in(0,T).\end{array}\right\}}$ (5)
For our case, $\alpha\in(0,1)$, $L$ and $T$ are positive real numbers. $f(x,t)$ is a given source term. The functions $f(x,t)$ and $\phi(x)$ are sufficiently smooth known functions.
In [37], the authors addressed a similar problem using the Atangana–Baleanu Caputo derivative by applying a direct linearization technique, which led to a reduced convergence order in the temporal direction. In contrast, we employ the Newton iteration method to achieve a higher convergence rate and improved accuracy. It is important to note that the scheme presented here is a novel contribution, as its application to solve problem (2) has not been previously explored in the literature.
@@
2 Derivation and Analysis of the Numerical Scheme
We assume the grid points as $t_{k}=k\tau,~k=0,1,\dots,N$ in time direction and for space direction, $x_{i}=ih,~i=0,1,\dots,M$, where $\tau=\frac{T}{N}$ and $h=\frac{L}{M},$ for $M,N\in\mathbb{Z}^{+}$. For the discretization of time-derivative of (2), we use finite difference method (FDM) in the definition (1). So at a grid point $(x_{i},t_{k})$, we have
@@ 2 Derivation and Analysis of the Numerical Scheme
$\approx$ $\sum_{j=0}^{k-1}\int^{t_{j+1}}_{t_{j}}(t_{k}-s)^{-\alpha}E_{\rho,1-\alpha}^{-\gamma}(\omega(t_{k}-s)^{\rho})u^{\prime}(x_{i},t_{j})ds$
$=$ $\frac{u(x_{i},t_{j+1})-u(x_{i},t_{j})}{\tau}\sum_{j=0}^{k-1}\int^{t_{j+1}}_{t_{j}}(t_{k}-s)^{-\alpha}E_{\rho,1-\alpha}^{-\gamma}(\omega(t_{k}-s)^{\rho})ds+r_{k}$
$=$ $\frac{1}{\mu}[a_{1}u_{i}^{k}-\sum_{j=1}^{k-1}(a_{k-j}-a_{k-j+1})u_{i}^{j}-a_{k}u_{i}^{0}]+r_{k},$ (6)
@@
$a_{n}=n^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{n})^{\rho})-(n-1)^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{n-1})^{\rho}),$ (7)
@@
Lemma 1.
[33] Let $u(.,t)\in C^{2}[0,1]$ then for any $\alpha\in(0,1)$, the truncation error $r_{k}$ occurred in approximating the CP derivative using FDM, given by
@@
$\lvert r_{k}\rvert$ $\leq C_{1}\frac{\tau^{2-\alpha}}{2}\max_{0\leq t\leq t_{k-1}}u^{\prime\prime}(x_{i},t_{k}),\$(8)
@@
We approximate the space derivatives of integer order from (2) as following,
$\frac{\partial^{2}u(x,t)}{\partial x^{2}}|_{(x_{i},t_{k})}$ $\approx\frac{u^{k}_{i+1}-2u^{k}_{i}+u^{k}_{i-1}}{h^{2}}:=\delta^{2}_{x}u_{i}^{k},$ (9)
$\frac{\partial u(x,t)}{\partial x}|_{(x_{i},t_{k})}$ $\approx\frac{u^{k}_{i+1}-u^{k}_{i-1}}{2h}:=\frac{1}{2h}\Delta u_{i}^{k},$ (10)
@@
$=\frac{1}{6h}(u_{i}^{k}\Delta u_{i}^{k}+\Delta(u_{i}^{k})^{2}).$ (11)
The discretizations (9) and (10) are simply central differences, while (11) is based on the Galerkin method using piecewise linear test functions [23, 30]. To get the fully discrete numerical scheme for the fractional Burgers equation (2), we substitute the approximations at point $(x_{i},t_{k})$ from (2), (9), and (11),
$\frac{1}{\mu}[a_{1}u_{i}^{k}-\sum_{j=1}^{k-1}(a_{k-j}-a_{k-j+1})u_{i}^{j}-a_{k}u_{i}^{0}]+\frac{1}{6h}(u_{i}^{k}\Delta u_{i}^{k}+\Delta(u_{i}^{k})^{2})=\delta^{2}_{x}u_{i}^{k}+f_{i}^{k},$ (12)
for $1\leq i\leq M,\ 1\leq k\leq N$, where $f_{i}^{k}=f(x_{i},t_{k})$. The initial and boundary conditions are as
$u^{k}_{0}=u^{k}_{M}=0,~~~~u^{0}_{i}=\phi(x_{i}).$ (13)
2.1 Stability and Convergence
First, we introduce some notations directly inspired by [23, 30], and we recommend these papers for a detailed explanation. Define $W=\{w=(w_{0},w_{1},\dots w_{M})|w_{0}=w_{M}=0\},$ and for any two elements $v,w\in W$ define inner product as $\langle v,w\rangle_{h}=\sum_{i=1}^{M-1}hv_{i}w_{i}$. The corresponding norms are $\|w\|_{h}=\sqrt{\langle w,w\rangle_{h}},~~~~\|w\|_{\infty}=\max_{1\leq i\leq{M-1}}|w_{i}|.$ Given the boundary conditions by (13), we have $\|w\|_{\infty}\leq\frac{\sqrt{L}}{2}\|\delta_{x}w\|_{h},$ where $\delta_{x}w_{i}=\frac{1}{h}(w_{i}-w_{i-i}).$
@@ 2.1 Stability and Convergence
$\langle\delta_{x}^{2}w_{1},w_{2}\rangle_{h}=-\langle\delta_{x}w_{1},\delta_{x}w_{2}\rangle_{h}~~~~\forall w_{1},w_{2}\in W,$ (14)
$\langle w\Delta w+\Delta(w)^{2},w\rangle_{h}=0,~~~~\forall w\in W,$ (15)
where $\delta_{x}^{2}$ and $\Delta$ are given by (9) and (10), respectively.
Lemma 2.
For any vector solution $U=(U_{0},U_{1},U_{2},\dots,U_{N})$ to (12)-(13), the following inequality is satisfied
@@
$-\frac{N^{1-\alpha}}{2}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N})^{\rho})U_{0}^{2},$ (16)
where $\{a_{n}|n=1,\dots,k\}$ are given by Eq. (7), and are decreasing function of $n$.
@@
By taking the left hand side of the inequality (2), we can simplify it as follows using simple well-known relations and inequalities
@@
$a_{N}=N^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N})^{\rho})-(N-1)^{1-\alpha}E_{\rho,2-\alpha}^{-\gamma}(\omega(t_{N-1})^{\rho})$, we have the inequality (2). ∎
Following theorem proves the boundedness of the numerical solution obtained from the proposed numerical scheme (12) which shows the scheme (12) is stable.
Theorem 3.
Let the solution of the proposed numerical scheme (12) with boundary and initial conditions (13) is $\{u_{j}^{k}|1\leq j\leq M-1,1\leq k\leq N\}$, then we have
@@
By taking inner product of Eq. (12) with $u^{k}$ then by summing up from 1 to $N$, we get
@@
In the last inequality above, we used Eq.(2) and Eq.(15). By the virtue of Young’s inequality: for all $a,b\in\mathbb{R}$ there exists $\varepsilon>0$ such that $ab\leq\varepsilon a^{2}+\frac{1}{4\varepsilon}b^{2},$ we have
@@
Now, we will discuss the convergence of our scheme for which we use energy method. Assume $v(x_{i},t_{k})=v^{k}_{i}$ be the exact solution of the problem (2)-(5), and let $\epsilon_{i}^{k}=v(x_{i},t_{k})-u^{k}_{i}$ be the error between exact solution and the approximated solution, then $\epsilon_{0}^{k}=\epsilon_{M}^{k}=0$ and $\epsilon_{i}^{0}=0~~\forall~~1\leq i\leq M-1,~~1\leq k\leq N$, and
@@
$-\frac{\mu}{6h}\left[\left(v_{i}^{k}\Delta\epsilon_{i}^{k}+\Delta(\epsilon_{i}^{k}v_{i}^{k})\right)+\epsilon_{i}^{k}\Delta v_{i}^{k}+\Delta(\epsilon_{i}^{k}v_{i}^{k})\right].$ (17)
@@
$+\mu\left[{(\gamma_{1})}_{i}^{k}-{(\gamma_{2})}_{i}^{k}-{(\gamma_{3})}_{i}^{k}-\frac{1}{6h}\left({(\gamma_{4})}_{i}^{k}-{(\gamma_{5})}_{i}^{k}\right)\right].$ (18)
We can easily calculate the following results by using initial and boundary conditions ( see [30, Lemma 3.6, Lemma 3.7])
$\langle{(\gamma_{4})}_{i}^{k},\epsilon^{k}\rangle_{h}=0,~~~~|\langle{(\gamma_{5})}_{i}^{k},\epsilon^{k}\rangle_{h}|\leq 3C_{0}h\|\epsilon^{k}\|_{h}^{2},$ (19)
$\sum_{k=1}^{N}\tau(\|{(\gamma_{1})}_{i}^{k}\|_{h}^{2}+\|{(\gamma_{2})}_{i}^{k}\|_{h}^{2}+\|{(\gamma_{3})}_{i}^{k}\|_{h}^{2})\leq C(\tau^{2-{\alpha}}-h^{2})^{2},$ (20)
@@
Now, in Eq. (18), taking product with $h\epsilon_{i}^{k}$ and then summing up for $i$ from 1 to M1M-1,
@@
Theorem 4.
The finite difference scheme (12) is convergent in the following sense
$\tau\sum_{k=1}^{N}\|\delta_{x}\epsilon^{k}\|_{h}^{2}\leq C(\tau^{2-\alpha}-h^{2})^{2},$ (21)
@@
3 Numerical Examples
Before conducting numerical examples we introduce Newton’s iterative method (see [30, 32]) to deal with the non-linear term in the numerical scheme (12).
3.1 Newton’s iterative method
Let us assume $\psi(u)=uu_{x}$, and hence the right hand side of (11) can be denoted as $\psi(u)^{k}_{i}$. By rewriting (12), we get